7x − y − z = 34
x − 3y + z = 6
x + 2y − z = 4
This is my first question! Find, x,y,z and make sure that they are all true!
4a − 3b = 1
6a − 8c = 1
2b − 4c = 0
This is my second question! find, x,y,z and make sure that they are all true!

Found 2 solutions by Edwin McCravy, jim_thompson5910:
Answer by Edwin McCravy(20055)

There are bunches of different ways to solve systems of equations.  
Which method are you studying?

7x −  y − z = 34
 x − 3y + z = 6
 x + 2y − z = 4

Here are 8 different solutions to that system.  If you substitute any one of
them, all three equations will be true.  

Solution 1:     x = 1   y = -8   z = -19
Solution 2:     x = 2   y = -6   z = -14
Solution 3:     x = 3   y = -4   z = -9
Solution 4:     x = 4   y = -2   z = -4
Solution 5:     x = 5   y = 0   z = 1
Solution 6:     x = 6   y = 2   z = 6
Solution 7:     x = 7   y = 4   z = 11
Solution 8:     x = 8   y = 6   z = 16

-------------------------------

4a − 3b = 1
6a − 8c = 1
2b − 4c = 0

Here's one solution to that system.  Substitute it and you'll get a true
equation:



Edwin


Answer by jim_thompson5910(35256)


I'll do the first problem to get you started

Given system of equations
7x − y − z = 34 ..... equation (1)
x − 3y + z = 6  ..... equation (2)
x + 2y − z = 4  ..... equation (3)


Solve equation (1) for z
7x - y - z = 34
7x - y - z+z = 34+z
7x - y = 34+z
7x - y - 34 = 34+z - 34
7x - y - 34 = z
z = 7x - y - 34 .... call this equation (4)

Plug this into equation (2). Simplify. Then solve for y.
x - 3y + z = 6
x - 3y + 7x - y - 34 = 6
8x - 4y - 34 = 6
8x - 4y = 40
2x - y = 10
y = 2x - 10 ... we'll use this later.

Move onto equation (3) and plug in z = 7x - y - 34. Simplify
x + 2y - z = 4
x + 2y - (7x - y - 34) = 4
x + 2y - 7x + y + 34 = 4
-6x + 3y + 34 = 4
-6x + 3y = -30
-2x + y = -10

Next replace y with 2x-10, which was from y = 2x-10 found earlier
-2x + y = -10
-2x + 2x - 10 = -10
-10 = -10
Both sides are the same value.
We have a consistent and dependent system.
There are infinitely many solutions


Go back to equation (4). Plug in y = 2x-10
z = 7x - y - 34
z = 7x - (2x - 10) - 34
z = 7x - 2x + 10 - 34
z = 5x - 24
We have now expressed z in terms of x

The variable y is in terms of x as well
y = 2x - 10

We can let x be the free variable. It is anything you want while y and z depend on what x is
x = any number
y = 2x-10
z = 5x-24

One way to express all the solutions is to write
(x, y, z) = (x, 2x-10, 5x-24)
This x,y,z triple will allow you to generate any solution you want

For instance, if x = 0, then
(x, y, z) = (x, 2x-10, 5x-24)
(x, y, z) = (0, 2*0-10, 5*0-24)
(x, y, z) = (0, -10, -24)
which is one solution

Or if x = 7, then
(x, y, z) = (x, 2x-10, 5x-24)
(x, y, z) = (7, 2*7-10, 5*7-24)
(x, y, z) = (7, 4, 11)
which is another solution.

Or if x = 5, then
(x, y, z) = (x, 2x-10, 5x-24)
(x, y, z) = (5, 2*5-10, 5*5-24)
(x, y, z) = (5, 0, 1)
which is another solution.

This process goes on forever because x can be any real number (and there are infinitely many of those).

note: x does not have to be the free variable. We could make either y or z have that role. I would ask your teacher which s/he prefers to have as the free variable. Often times, it is sufficient enough to say "infinitely many solutions" and not have to specify the solution form.